OpenStudy (anonymous):

Solve for w in P = 2w + 2l, if P = 38 and l = 12.

OpenStudy (whpalmer4):

\[P = 2w + 2l\]Subtract \(2l\) from both sides \[P - 2l = 2w\]Divide both sides by 2 \[\frac{P-2l}{2} = w\]Now plug in your numbers to find the value of \(w\)

OpenStudy (anonymous):

I still don't really get it though. .-.

OpenStudy (whpalmer4):

What don't you get? The algebra, or how to plug in your numbers?

OpenStudy (anonymous):

how to plug in the numbers.

OpenStudy (whpalmer4):

Okay, well, where you see "P" replace it with "(38)". Where you see "l", replace it with "(12)". What do you get?

OpenStudy (anonymous):

38/12 is 3?

OpenStudy (whpalmer4):

Uh, no. \[w = \frac{P-2l}{2} = \frac{(38)-2(12)}{2}\] What is 2*12? What is 38 - 2*12? What is (38-2*12)/2?

OpenStudy (anonymous):

2x12 is 24 and 38-24 is 14 and 14/2 is 7

OpenStudy (anonymous):

oooh okay. I get it. Thank you it helped a lot.

OpenStudy (whpalmer4):

You got it!